OpenStudy (bigbosssaint21):

Choose the correct simplification of (5xy^7)2(y^2)3. A: 10x^2y^20 B: 25x^2y^20 C: 25x^2y^14 D: 10x^2y^14

OpenStudy (anonymous):

I'll assume it's \[\Large (5xy^7)^2(y^2)^3.\] correct?

OpenStudy (bigbosssaint21):

lol yea, sorry

OpenStudy (anonymous):

It's ok, we you know the rules of exponents or no?

OpenStudy (bigbosssaint21):

um, i could prob use a refresher

OpenStudy (anonymous):

Ok, well basically when you have bases and exponents 

OpenStudy (anonymous):

When you have like bases multiplied together, you add the exponents so if you had something like \[\Large x^2 \times x^3=x^{2+3}=x^5\]

OpenStudy (anonymous):

If you have a base/exponent raised to a power, you multiply \[\Large (x^2)^4=x^{2 \times 4}=x^8\]

OpenStudy (anonymous):

Now when you have an expression (a bunch of unlike bases and their exponents multiplied together) and they're raised to a power, you have to raise each different base to that power \[\Large (5x^2y^3)^3=5^3x^6y^9\] Do you follow me so far @Bigbosssaint21

OpenStudy (bigbosssaint21):

i believe so

OpenStudy (anonymous):

Ok, now we're going to apply this to the problem \[\Large (5xy^7)^2(y^2)^3=?\]

OpenStudy (anonymous):

Can you go through and tell me what they would come out to? Rermeber that the 2 is going to go the 5, x and y^7 and then that will be multiplied by the y^6 which is just the (y^2)^3 2x3=6

OpenStudy (bigbosssaint21):

(5^23x^2y^16)?

OpenStudy (anonymous):

If you check the equation button below the comment bar, it'll help make your equations easier to understand so it'd be \[\Large (5^2x^2y^{7^{2}}) \times y^6\]

OpenStudy (bigbosssaint21):

oh ok

OpenStudy (anonymous):

Yes and the first y will be y^14 and then it will add to the y^6 (li0ke bases) and give you 25x^2y^20

OpenStudy (bigbosssaint21):

ok, thank you =)